Presented below are data relating to labor for Verde Appliance Repair Shop. Repair-technicians wages Fringe benefits Overhead $119,228 34,112 51,332 The and desired profit margin per hour is $16.54. The material loading charge is 60% of invoice cost. Verde estimates that 4,100 labor hours will be worked next year. If Verde repairs a dishwasher that takes 1.3 hours to repair uses parts of $72.53, compute the bill for the job. (Round answer to 2 decimal places, e.g. 10.50.) Total cost
Answer #1

Let us first calculate the expenses per labor hour.

Repair-technicians' wages per labor hour = $119,228 / 4,100 labor hours = $29.08

Fringe benefits per labor hour = $34,112 / 4,100 labor hours = $8.32

Overhead per labor hour = $51,332 / 4,100 labor hours = $12.52

Calculation of bill amount for the repair of dishwasher :

Repair-technicians' wages

[ $29.08 * 1.3 labor hour ]

$37.804

Fringe benefits

[ $8.32 * 1.3 labor hour ]

$10.816

Overhead

[ $12.52 * 1.3 labor hour ]

$16.276
Material parts used $72.53

Materials loading charge

[ 60% * $72.53 ]

43.518

(+) Profit

[ $16.54 * 1.3 labor hour ]

$21.502
Bill amount $202.45